OpenStudy (anonymous):

I am stuck on this homework problem. All solutions in [0, 2π) for sin2+√3cosx=0. That is how it is written, but is it correctly grouped as: sin(2+√3)cos(x)=0 ? We are using double-angle identities.

OpenStudy (anonymous):

Typo mistake. It must be sin2x. \[\huge \sin2x+\sqrt{3}\cos x=0\] \[\huge 2\sin x\cos x+\sqrt{3}\cos x=0\] \[\huge \cos x(2\sin x+\sqrt{3})=0\] \[\huge \cos x=0\] OR \[\huge \sin x=-\frac{\sqrt{3}}{2}\] THEREFORE: x=\frac{\pi}{2}, \frac{3\pi}{2}, \frac{4\pi}{3}, \frac{5\pi}{3}

OpenStudy (anonymous):

\[\huge x=\frac{\pi}{2}, \frac{3\pi}{2}, \frac{4\pi}{3}, \frac{5\pi}{3}\]

OpenStudy (anonymous):

Sorry about the last line in the first post. Rewrote it in my second post.

OpenStudy (anonymous):

Thank you Azteck. The grouping confused me and I overlooked factoring the cos.